OpenStudy (anonymous):

Complete the factoring 15x^2-6x= 3x( )

OpenStudy (anonymous):

5x-2

OpenStudy (anonymous):

Another fun method: 3x*a = 15x^2 a = 15x^2/3x = 5x 3x*b = -6x b = -6x/3x =-2 (5x-2)